OpenStudy (anonymous):

Write a word problem that can be solved by working backward. Include at least two steps in the problem. Then, show the answer and how the problem can be solved by working backward.

OpenStudy (mathstudent55):

Think of a number. Then do several operations to the number and find the result. Once you know the result, word the problem as follows. What is the number that when [insert here all the operations you did in order] the result is [the result of your calculation].

OpenStudy (anonymous):

umm i dont get it

OpenStudy (mathstudent55):

You need to create your own problem. You need to start with the solution. Let's make the solution 2.

OpenStudy (mathstudent55):

Now we come up with several operations to do starting with 2. 1. Add 8 2 + 8 = 10 2. Multiply by 4 10 * 4 = 40 3. Divide by 5 40/5 = 8

OpenStudy (mathstudent55):

Now you word the problem based on the operations you did. You don't mention the number you started with. You just give the final result of all the operations and the person has to work the operations backwards to find the number.

OpenStudy (mathstudent55):

For example, you can state: I picked a number, added 8 to it, then multiplied by 4, then divided by 5, and the result was 8. Find the number.

OpenStudy (mathstudent55):

We know the answer is 2 because we made up the problem. The way for someone to find the answer would be to start with the final result, 8. Then do all the operations mentioned in reverse and in reverse order.

OpenStudy (anonymous):

@giana_leee the previous problem you just posted is the example fo the same..

OpenStudy (mathstudent55):

Start with 8. To undo the last operation of divide by 5, we multiply by 5. 8 * 5 = 40

OpenStudy (mathstudent55):

To undo the multiply by 4, we divide by 4. 40/4 = 10

OpenStudy (anonymous):

all this is cofusing

OpenStudy (mathstudent55):

To undo the add 8, we subtract 8. 10 - 8 = 2 There is the answer. The number picked was 2.

OpenStudy (mathstudent55):

Go over it a few times. Maybe then it'll make more sense.

OpenStudy (anonymous):

alright